determine the number of terms n in each geometric series.
Given that a1=3 r=6 and Sn=129

Answer by greenestamps(13334)



a1=3; r=6....

S1 = 3

S2 = 3+18 = 21

S3 = 3+18+108 = 129

ANSWER: n=3